# Large Language Models Could Be Rote Learners

### v4

@@ 3.3 Measuring LLM Capability with TrinEval
To understand how LLMs store and manipulate knowledge, Allen-Zhu and Li (2023) created a fictional biography dataset that enumerates various attributes (e.g.,, names, jobs, universities) and trained LLMs on this dataset. They employed a linear query-based probing method to uncover correlations between the entity token embeddings and the associated attributes, revealing that where LLMs encode knowledge, e.g., under person names or sequence of the knowledge mention, is crucial for robust mastery of knowledge. This insight leads us to believe that entity tokens, which should ideally store related knowledge, are the target for evaluating an LLM’s genuine capability.
@@
However, applying this method to real-world datasets, such as MMLU, presents challenges. Unlike controlled datasets with explicitly defined attributes, real-world data includes a far broader range of possible knowledge. As a result, we cannot enumerate all potential attributes and directly apply linear probing. To this end, we propose TrinEval, a verbal query probing method that reformulates MCQs around a knowledge-centric trinity: knowledge entity, attribute, and context. TrinEval is a pluggable augmentation on any MCQ benchmarks and could expose the genuine capability of LLMs by verifying whether they have correctly encoded knowledge. We next explain the elements in the trinity and how to reformulate.
@@
Knowledge entity. We suppose that if an LLM has mastered some knowledge, the key information pertinent to the knowledge should be encoded within a few subject tokens, namely knowledge entity, to support efficient retrieval. By isolating these tokens, TrinEval ensures that only the essential information is considered.
@@
Attribute. The attribute acts as a verbal probe to guide the model focusing on the specific feature or property of the knowledge entity being inquired. This mechanism allows TrinEval to isolate and assess the model’s understanding of the critical aspects of the questioning subject.
@@
Context. In a certain portion of questions, the conditions or background context can significantly influence the solution approach. By explicitly including context in the evaluation process, TrinEval helps the model account for relevant situational details that might otherwise be overlooked, ensuring that the model’s answer is based on a comprehensive understanding of the problem.
@@
By extracting the core and necessary question information in this trinity format, the reformulation by TrinEval is knowledge-preserving for the purposes of assessment. In the meanwhile, it completely destructs the original token sequence, effectively reducing the influence of memorization. We will empirically verify these properties through experiments. The reformulation is completed by a two-round reflection-based prompting method, with detailed procedure (Alg. 1) and related prompts available in Appendix B. Given an MCQ $x=(x_{Q},x_{O},x_{W})$, it first queries a capable reformulation LLM to derive the knowledge entity $x_{E}$, attribute $x_{A}$, and context $x_{C}$ from the original $x$. The LLM is instructed that the triplet should be sufficient for answering the question correctly, without including the answer option itself, ensuring the integrity of the evaluation. The same LLM then assesses whether the triplet contains all necessary information and no redundant information (typically, the rote-memorization), in the meanwhile, yields a rationale $x_{L}$ as reflection (Shinn et al., 2024; Yao et al., 2022). If it does, the triplet is returned as the re-formulated question. Otherwise, the reformulation model refines the extraction, taking as input $x_{E}$, $x_{A}$, $x_{C}$, and $x_{L}$, and re-evaluates the updated triplet.
Finally, prompting with the extracted $x_{E}$, $x_{A}$, and $x_{C}$ as well as options $x_{O}$, we inspect the generation probability of the ground-truth answer $x_{W}$ as the first token to measure capability:
$F_{c}(x,G)=p_{G}(x_{W}|x_{E},x_{A},x_{C},x_{O}).$ (2)
As can be seen, the $F_{c}$ metric retains the necessary knowledge-centric information while discarding unnecessary biases, especially the rote memorization of LLMs, which leads to the quantification of genuine capability of LLMs.
4 Experiments
Q3. What insights does TrinEval provide into the distinction between rote memorization and genuine capability in LLMs?
4.1 Experiment Setup
Models. We employ API-accessible commercial LLMs for question reformulation in TrinEval, specifically gpt-4o-2024-08-06 (GPT) (OpenAI, 2024) and qwen-max-2024-09-19 (Qwen) (Yang et al., 2024; Team, 2024). For model evaluation, we utilize open-source LLMs to obtain token-level logits. Our experiments focus on three widely adopted models: Llama2-7B (Llama) (Touvron et al., 2023), Mistral-7B-v0.2 (Mistral) (Jiang et al., 2023), and Vicuna-v1.5-7B (Vicuna) (Zheng et al., 2023b). These models are accessed from Hugging Face and implemented with the transformers library, allowing us to compute the log-probabilities of generated tokens for fine-grained analysis. All evaluations are conducted on a single NVIDIA A100 GPU with 40GB of memory. Default generation parameters are used, and greedy decoding is applied throughout to ensure reproducibility.
Benchmarks. We evaluate LLMs on the widely used MMLU benchmark (Hendrycks et al., 2020), in light of growing evidence of data contamination (Touvron et al., 2023) [^3] https://hitz-zentroa.github.io/lm-contamination/ in many recent LLMs on this data. MMLU spans 57 subjects across STEM, humanities, social sciences, and others, providing a comprehensive assessment of model capabilities. We find duplicated MCQs across subjects on MMLU, and thus de-duplicate the dataset, resulting in a test set of 14,006 unique questions.
Evaluation. With commercial LLMs, we evaluate model performance by extracting predicted answers with regular expressions. For open-source models, we access the output probability of the first generated token (i.e., corresponding to multiple-choice option IDs A/B/C/D) to compute quantitative performance metrics.
4.2 Q1. Is TrinEval Knowledge-preserving?
Figure 3: Knowledge-preserving validation of TrinEval reformulation. The green, blue, and overlapping regions represent the sets of MCQs correctly answered in the original format, TrinEval format, and both formats, respectively. Best viewed in color.
We begin by verify whether TrinEval’s reformulation preserves essential knowledge, a prerequisite for reliable knowledge assessment. Specifically, our goal is to ensure that (1) the reformulation does not omit critical information that would cause originally correctly answered questions to be answered incorrectly, and (2) it does not introduce anomalous or unexpected content that could artificially inflate performance.
Following the complete TrinEval reformulation pipeline, we obtain 4,343 qualified MCQs with corresponding knowledge entities, attributes, and contexts using Qwen, and 4,645 MCQs with associated triplets using GPT. We then prompt both Qwen and GPT to answer their respective MCQs in both the original and reformulated triplet-based formats. Example MCQs and the questioning prompt templates in both formats are provided in Table 2 in the Appendix; results are summarized in Fig. 3.
Among the qualified MCQs reformulated by Qwen and GPT, 4,133 and 4,409 are answered correctly in at least one of the two formats, respectively. Notably, the majority of these are answered correctly in both the original and TrinEval formats, accounting for 90.3% with Qwen and 90.4% with GPT, indicating strong consistency in terms of problem-solving across formats. Only 121 questions for Qwen (2.9%) and 226 for GPT (5.1%) are answered correctly exclusively in the original format, suggesting that the TrinEval reformulation does not omit essential information. Conversely, 278 questions for Qwen (6.7%) and 197 for GPT (4.5%) are answered correctly only in the reformulated format, which is comparable in general to the numbers for the original format alone. This further suggests that the reformulation does not introduce extra information that might artificially enhance model performance. Collectively, these findings provide strong evidence that TrinEval preserves the core knowledge necessary for accurate answering, satisfying the requirements for reliable capability evaluation.
4.3 Q2. Can TrinEval Reduce Memorization?
(a) Evoc. with Llama2 and dev-fsp.
(b) Evoc. with Mistral and dev-fsp.
(c) Evoc. with Vicuna and dev-fsp.
(d) Evoc. with Llama2 and seq-fsp.
(e) Evoc. with Mistral and seq-fsp.
(f) Evoc. with Vicuna and seq-fsp.
Figure 4: The results of memorization evocation (evoc.) under various dataset-related context, with green and blue curves referring to the memorization difference $\Delta F_{m}$ in the original and TrinEval formats, respectively. In the x-axis, ‘clean’, ‘meta’, ‘dev-fsp’, and ‘seq-fsp’ stand for without dataset-related context, with the name of the dataset, with few-shot prompt from the development set, and with few-shot prompt from the test set ahead of the testing question. These results of $\Delta F_{m}$ indicate the growing memorization effect given the increasing dataset-related information in general. However, the $\Delta F_{m}$ by TrinEval under the strongest memory evocation context remains consistently lower than the one in the original format, e.g., the red dashed line.
In this subsection, we investigate whether the proposed TrinEval reformulation can suppress unnecessary memorization in LLMs, thereby isolating and revealing the models’ genuine capabilities under various circumstances. Following the methodology of Bordt et al. (2024), we introduce dataset-specific cues into the questioning prompts and assess to what extent TrinEval mitigates memorization elicited by such prompts.
To evoke memorization, we embed dataset-related metadata into the input prompts, i.e., the dataset name and in-context few-shot examples drawn from the same dataset (see Appendix E). For this set of experiments, we focus on LLaMA, Mistral, and Vicuna, as their open-source implementations allow access to token-level output probabilities, enabling us to compute the memorization metric $F_{m}$. Since $F_{m}$ lacks a defined absolute zero point indicating complete absence of memorization, we use the $F_{m}$ value obtained from the original MCQ format without any additional context (i.e.,, the vanilla MCQ) as a reference baseline. We then visualize the average change in $F_{m}$ for each memorization-evoking method relative to this baseline.
Specifically, we hypothesize that unintentional data contamination may occur when LLMs are pretrained on datasets scraped from public sources (e.g.,, Hugging Face), where nearby examples from the same dataset may be concatenated during corpus construction. Following the pretraining mechanism of next-token prediction and prior findings from Carlini et al. (2022a), we posit that the inclusion of dataset-similar samples in the prompt may inadvertently trigger memorization. Accordingly, we design a sequence of memorization-evoking perturbations by progressively increasing the contextual cues: (1) providing only the dataset name, (2) adding few-shot examples from the same dataset (i.e.,, samples in the development set or ahead examples in the test set of the same subject of MMLU), and (3) combining both types of context. For each MCQ, we compute the change in $F_{m}$ relative to the reference baseline, capturing the degree to which memorization is elicited by specific context in original or TrinEval formats. Results are illustrated in Fig. 4.
Consistent with the observations by Bordt et al. (2024), the results show that $F_{m}$ increases as more dataset-specific context is introduced, indicating stronger memorization effects. However, across all three tested LLMs, the $\Delta F_{m}$ curve for TrinEval remains consistently lower than that of the original MCQ format (i.e., the red dashed line) regardless of various context used. Notably, even under the strongest memorization-evoking setting, TrinEval’s absolute $F_{m}$ still stays below the vanilla baseline. These findings provide compelling evidence that TrinEval significantly mitigates the influence of memorization, effectively disentangling spurious recall from genuine model capability.
4.4 Q3. TrinEval’s Findings on LLM Memorization and Genuine Capability
(a) $F_{m}$ v.s. $F_{c}$ with Llama on Qwen-extracted triplets.
(b) $F_{m}$ v.s. $F_{c}$ with Mistral on Qwen-extracted triplets.
(c) $F_{m}$ v.s. $F_{c}$ with Vicuna on Qwen-extracted triplets.
(d) $F_{m}$ v.s. $F_{c}$ with Llama on GPT-extracted triplets.
(e) $F_{m}$ v.s. $F_{c}$ with Mistral on GPT-extracted triplets.
(f) $F_{m}$ v.s. $F_{c}$ with Vicuna on GPT-extracted triplets.
Figure 5: The distribution of MCQs based on memorization metric $F_{m}$ vs. capability metric $F_{c}$. According to the values of $F_{m}$ and $F_{c}$, we separate the MCQs into 25 groups and visualize the MCQ distribution from weak to strong with heatmaps.
In this subsection, we aim to explicitly examine the relationship between memorization and genuine problem-solving capability in LLMs, using the metrics $F_{m}$ (memorization) and $F_{c}$ (capability). Since commercial API-based models do not expose full vocabulary-level output probabilities, we focus on open-source LLMs for computing these metrics. After computing the $F_{m}$ and $F_{c}$ scores for all qualified MCQs in TrinEval format, we divide the values of $F_{m}$ and $F_{c}$ into five equal intervals, respectively, resulting in 25 distinct groups of MCQ samples based on their joint distribution. We then use heatmaps to visualize the distribution of samples across these groups, thereby revealing the relationship between memorization and capability for each tested model.
As shown in Fig.5, the majority of MCQs cluster in the lower-left and upper-right corners of the heatmap. This pattern is evident and consistent across our tests. For instance, in the results of LLaMA using Qwen-extracted triplets, 38.57% of MCQs fall within the combined lower-left and upper-right $2\times 2$ grid regions, yielding a Pearson correlation of -0.7755 (p-value $<0.05$) for $F_{m}$ vs. $F_{c}$ of examples within the two square regions; expanding to the $3\times 3$ regions increases coverage to 74.17%, with a stronger correlation of -0.8124 (p-value $<0.05$). Similarly, for Mistral on Qwen triplets, 44.90% of MCQs lie in the two $2\times 2$ corner regions with a correlation of -0.8722 (p-value $<0.05$), and 80.82% in the two $3\times 3$ corner regions with a correlation of -0.8794 (p-value $<0.05$). We take the MCQs within the upper-right 2 X 2 squares in Fig. 5 as the ones memorized by rote and compute the ratio of 19.6% by averaging the percentage of all evaluated LLMs. Similarly, 20.5% is computed as the ratio of MCQs within the lower-left 2 X 3 squares with genuine capability. Additional statistics results are reported in Table 3 in Appendix. These findings consistently indicate a negative correlation between memorization and capability: MCQs with lower memorization scores are more likely to reflect genuine problem-solving capability, whereas high memorization levels correspond to degraded performance, suggesting overfitting or shallow recall.
To interpret these results, we draw an analogy between LLM behavior and the human memory system, which comprises two key components: Short-Term Memory (STM) and Long-Term Memory (LTM) (Shiffrin, 2003). Neurobiological studies reveal that STM relies on transient synaptic protein synthesis with limited temporal persistence and functional scalability. In contrast, LTM is constructed through stabilized neuronal memory traces that constitute an enduring knowledge framework. This neural architecture not only supports STM operations as a cognitive substrate but also enables sophisticated information generalization across diverse contexts. This dichotomy aligns with recent observations in LLMs. As shown by Allen-Zhu and Li (2023) and Ovadia et al. (2023), models trained on corpora with diverse rephrasings exhibit stronger generalization than those trained solely on original formulations. When trained on a single fixed corpus format, LLMs, due to their strong memorization capacity, tend to activate their STM system to memorize at the token level, capturing surface patterns rather than abstract knowledge. In this sense, LLMs are potential rote learners. Reformulations like TrinEval appear to facilitate the encoding of knowledge in a more principled, structured and reusable form, akin to LTM, thereby supporting more generalized and robust evaluation. Further experimental details and results are provided in Appendix C.
Figure 6: Averaged distance of each MCQs between the closest $1\%$ MCQs’ embeddings. ‘Rote Memorization’ refers to MCQ within lower left $2\times 2$ squares that typically exhibits high memorization metric $F_{m}$ and low capability $F_{c}$ while the ‘Genuine Capability Learning’ stands for MCQ lies within the upper right $2\times 2$ squares with low $F_{m}$ but high $F_{c}$. Further results are shown in Appendix F.
To further validate our hypothesis, we analyze the semantic proximity of MCQs within the qualified MMLU dataset by computing their embeddings and measuring the average distance to their closest 1% of samples. Our underlying assumption is that LLMs could better master a knowledge point if it is expressed in various formats in the training corpus. As a result, these knowledge ‘rephrases’ are also encoded densely via LLMs. By selecting only the top 1% nearest neighbors, we aim to retain the samples that are semantically-correlated in knowledge description. We then compare the mean embedding distances of MCQs located in the lower-left (strong memorization, weak capability) and upper-right (weak memorization, strong capability) $2\times 2$ regions of the heatmap in Fig. 5. The corresponding results are presented in Fig. 6. Interestingly, we observe that the average embedding distance among the Genuine Capability Learning MCQs is much lower than (nearly half) that of the Rote Memorization MCQs. This suggests that memorized MCQs are more sparsely distributed in the embedding space, while non-memorized, capability-driven MCQs tend to form tighter semantic clusters. This observation also aligns well with the cognitive analogy of STM and LTM: rote memorization leads to fragmented, context-specific encoding, whereas genuine capability emerges from structured, reusable representations. The results of alternative thresholds are shown in Table 5. Though it is well believed that memorization may lead to better but cheating performance of LLMs, we prove that the more LLMs memorize, the worse they are at solving problems.
5 Conclusion
This study provided a novel perspective on benchmark contamination in LLM evaluation, reframing it as an inherent aspect of learning. This perspective led us to explore the relationship between memorization and genuine capability in LLMs. Through our empirical investigation, we observed a surprising result: LLMs performed worse on memorized MCQs compared to those not, suggesting that superficial memorization may undermine problem-solving ability rather than enhance it. This finding also implies the existence of two distinct learning paradigms in LLMs: rote memorization and genuine capability learning.
To disentangle them, we proposed TrinEval, a novel evaluation method that reformulates MCQs into a knowledge-centric trinity, thus separating the influence of memorization from genuine knowledge application. Experiments validated both the knowledge-preserving and memorization-reducing properties of this approach. Based on that, TrinEval reveals the in-robustness of LLMs’ knowledge learning, e.g., popular open-source LLMs memorize 20.5% of knowledge points by rote without understanding in MMLU. As such, we believe this work lays the groundwork for future studies on improving LLM knowledge robustness and more thorough evaluation.
6 Limitations
Our limitations are mainly two points. First, though our proposed TrinEval retrains the problem-solving ability of the LLMs and obtains stronger robustness, it is not a dynamical re-organizing method that can still be leaked and pre-experienced during training. On the one hand, we appeal to the LLM developers not to use this re-organizing method as part of the training corpus. On the other hand, future works will be focused on developing dynamic evaluation method (Zhu et al., 2023, 2024). Second, we did not give a clear exploration on how and why the more LLMs memorize, the less the capability of the LLMs obtains. In future work, we will also look into the mechanism of the training and structure of LLMs for a thorough study of the phenomenon.
In this section, we introduce the details of the processed dataset and the prompts for extraction. MCQs from some subjects contain similar or identical options [^4] E.g., the options of MCQs in the subject, moral_scenarios, are all identical (‘Wrong, Wrong’, ‘Wrong, Not wrong’, ‘Not wrong, Wrong’ and ‘Not wrong, Not wrong’).. With the provided 5-shot prompt, options of MCQs from these subjects can be easily extracted, leading to a high False-positive ratio. In order to avoid the influence of the few-shot prompt on the option extraction, we eliminate MCQs in which any of the options have appeared twice in the dataset. After deduplication, we obtain 14,006 MCQs for evaluation. The extraction prompt and the detailed statistics are shown in the following text and Tab. 1.
Prompt template for extraction: Prompt template for pre-investigation on LLM Memorization w.r.t. Capability:
You are an expert of multiple choice questions of MMLU dataset. The following are multiple-choice questions (with answers) about [subject]. [examples] [question] Options: A. You are an expert of multiple choice questions of MMLU dataset. The following are multiple choice questions (with answers) about [subject]. [examples] [question] Options: A. [content for option A] B. [content for option B] C. [content for option C] D. [content for option D] Answer:
{tblr}
cells = c, cell21 = r=3, cell51 = r=3, cell81 = r=3, vline2-3 = 1-2,5,8, vline2-3 = 3-4,6-7,9-10, hline1-2,5,8,11 = -, Model & Subset Simple Pro MMLU
Llama memorized 912 70 982
non-mem. 6,548 6,476 13,024
all 7,460 6,546 14,006
Mistral memorized 879 36 915
non-mem. 6,581 6,510 13,091
all 7,460 6,546 14,006
Vicuna memorized 893 16 909
non-mem. 6,567 6,530 13,097
all 7,460 6,546 14,006
Table 1: Statistics of memorized and non-memorized questions by Llama2-7B, Mistral-7B-v0.2, and Vicuna-v1.5-7B in MMLU.
Appendix B Details of TrinEval
In this section, we introduce the details of the proposed TrinEval. The pseudo-code is shown in the Alg. 1. The prompts used are also shown below. Note that the potential data leakage is often caused by the data crawled on the Huggingface dataset site. Thus, we also provide the original text of MCQs in the format on the Huggingface dataset site to mimic the data contamination with in-context learning.
@@ Appendix B Details of TrinEval
13:        Discard the MCQ, return $None$;

### v5

@@ III-B Measuring LLM Capability with TrinEval
To understand how LLMs store and manipulate knowledge, Allen et al. [18] created a fictional biography dataset that enumerates various attributes (e.g.,, names, jobs, universities) and trained LLMs on it. They employed a prompt-based linear probing method to uncover correlations between the entity token embeddings and the associated attributes, revealing that where LLMs encode/store knowledge, e.g., under person names or a sequence of knowledge augmentation, is crucial for robust mastery of knowledge. This insight leads us to believe that entity tokens, which should ideally store knowledge, are the target for evaluating an LLM’s genuine capability.
@@
However, applying this method to real-world datasets, such as MMLU, presents challenges. Unlike controlled datasets with explicitly defined attributes, real-world data includes a far broader range of possible knowledge. As a result, we cannot enumerate all potential attributes and directly apply linear probing. To this end, we propose TrinEval, a verbal query probing method that reformulates MCQs around a knowledge-centric trinity: knowledge entity, attribute, and context. TrinEval is a pluggable augmentation on any MCQ benchmarks and could expose the genuine capability of LLMs by verifying whether they have correctly encoded knowledge. In contrast to Allen et al. [18], TrinEval introduces a verbal query-based probing mechanism. It reformulates MCQs into a structured triplet format (entity–attribute–context), enabling direct assessment of knowledge encoded in entity tokens—without relying on model internals. This shift from embedding-space analysis to verbalized knowledge extraction is a conceptual and methodological advance. Generally, we instruct LLMs to extract the triplet according to the prompt and check and refine it if they are not qualified. This systematic validation is absent in most prior rephrasing methods and is empirically shown to be knowledge-preserving without introducing extraneous cues. The reformulation is completed by a two-round reflection-based prompting method, with the detailed procedure shown in Alg. 1 and related prompts available in Fig. 2 in Appendix B. Here, we explain the elements of the trinity and how to reformulate them.
@@
Knowledge entity. We suppose that if an LLM has mastered some knowledge, the key information pertinent to the knowledge should be encoded within a few subject tokens, namely, knowledge entity, to support efficient retrieval. By isolating these tokens, TrinEval ensures that only the essential information is considered.
@@
Attribute. The attribute acts as a verbal probe to guide the model to focus on the specific feature or property of the knowledge entity being inquired about. This mechanism allows TrinEval to isolate and assess the model’s understanding of the critical aspects of the questioning subject.
@@
Context. For some questions, conditions or background contexts can significantly influence the solution approach. By explicitly including context in the evaluation process, TrinEval helps the model account for relevant situational details that might otherwise be overlooked, ensuring that the answer is based on a comprehensive understanding of the problem.
Traditional rephrasing often relies on lexical variations (e.g., synonym replacement or sentence restructuring), which may still preserve memorization cues. In contrast, by extracting the core and necessary question information in this trinity format, the reformulation by TrinEval is knowledge-preserving for the purposes of assessment. In the meantime, it completely destroys the original token sequence, effectively reducing the influence of memorization. We will empirically verify these properties through experiments.
@@
Based on the framework, we locate which token the knowledge is stored within and expose the genuine capability of LLMs with the trinity. Given an MCQ $x=(x_{Q},x_{O},x_{W})$, it first queries a capable reformulation LLM to derive the knowledge entity $x_{E}$, attribute $x_{A}$, and context $x_{C}$ from the original $x$. The LLM is instructed that the triplet should be sufficient for answering the question correctly, without including the answer option itself, ensuring the integrity of the evaluation. In order to mimic the data contamination with in-context learning, we also present the original MCQ text in the format on the Huggingface dataset site as the “**Fact:**” part in the prompt, given that the potential data leakage is often caused by the data crawled on the Huggingface dataset site. The same LLM then assesses whether the triplet contains all necessary information and no redundant information (typically, the rote-memorization), meanwhile, yields a rationale $x_{L}$ as reflection [42, 41]. If it does, the triplet is returned as the re-formulated question. Otherwise, the reformulation model refines the extraction, taking as input $x_{E}$, $x_{A}$, $x_{C}$, and $x_{L}$, and re-evaluates the updated triplet. The details of the reorganization algorithm are shown as follows. Reflection prompts are shown in Appendix B.
@@
13:   Discard the MCQ, return NoneNone;